OpenStudy (anonymous):

Solve the system by the method of substitution. y = 2x + 3 y = 3x + 1

OpenStudy (lopus):

is easy y=2x+3 === (3x + 1)=2x+3

OpenStudy (anonymous):

just distribute

OpenStudy (ybarrap):

just subtract

OpenStudy (texaschic101):

no....solve for x by getting the x's on one side and everything else on the other side

OpenStudy (texaschic101):

do not distribute

OpenStudy (anonymous):

o kk

OpenStudy (ybarrap):

\[(2x+3)-(3x+1)=0\]

OpenStudy (texaschic101):

that will work as well

OpenStudy (anonymous):

thx

OpenStudy (texaschic101):

thats what we are here for :)

OpenStudy (anonymous):

lol you guys are really nice

OpenStudy (anonymous):

i got x=-2

OpenStudy (anonymous):

how do i find y

OpenStudy (texaschic101):

3x + 1 = 2x + 3 3x - 2x = 3 - 1 x = 2 now sub 2 in for x in either of the original equations to find y

OpenStudy (anonymous):

o kk

OpenStudy (anonymous):

sry

OpenStudy (anonymous):

y=7

OpenStudy (texaschic101):

you got it...way to go :)

OpenStudy (anonymous):

thx again

OpenStudy (texaschic101):

no problem :)